# MCP-Zero: Active Tool Discovery for Autonomous LLM Agents

### v3

@@ Abstract
Current LLM agents inject thousands of tool schemas into prompts, creating massive context overhead and reducing models to passive tool selectors rather than autonomous agents. We introduce MCP-Zero, an active agent framework that restores tool discovery autonomy to LLMs themselves. Instead of overwhelming models with all available tools, MCP-Zero enables agents to actively identify capability gaps, and request specific tools on-demand, transforming them from large-scale retrievers into genuine autonomous agents. The framework operates through three core mechanisms: (1) Active Tool Request, where models autonomously generate structured requests specifying their exact tool requirements; (2) Hierarchical Semantic Routing, a two-stage algorithm that matches requests to relevant servers and tools through improved semantic alignment; (3) Iterative Capability Extension, enabling agents to progressively build cross-domain toolchains while maintaining minimal context footprint. We also construct MCP-tools, a comprehensive dataset of 308 MCP servers and 2,797 tools from the official Model-Context-Protocol repository. Experiments demonstrate that MCP-Zero preserves agent autonomy while achieving substantial efficiency gains: (i) accurate tool selection from nearly 3k candidates across 248.1k tokens; (ii) 98% reduction in token consumption on APIBank while maintaining high accuracy; and (iii) consistent multi-turn performance that scales with tool ecosystem growth. This work establishes active tool discovery as a fundamental design pattern for scalable autonomous agent systems. The code and dataset is released at https://github.com/xfey/MCP-Zero.
@@ 1 Introduction
Autonomous agents represent a fundamental shift in AI system design: from passive text processors to active decision-makers capable of perceiving their environment, identifying capability gaps, and taking purposeful actions [30, 8, 1]. The emergence of function-calling mechanisms has enabled large language models (LLMs) to transcend their parametric boundaries, leveraging external tools, APIs, and execution environments to accomplish complex real-world tasks [21, 26, 14, 18].
@@ 4.1 Dataset Construction
Structured Information Extraction. We employed Qwen2.5-72B-Instruct [24] with carefully designed few-shot examples to extract structured information following a standardized schema:
@@
The server_summary field represents a key innovation: LLM-generated comprehensive summaries that distill server capabilities while excluding operational details. This design enables effective semantic matching across diverse query formulations, from formal API specifications to natural language capability descriptions.

### v4

@@ Abstract
True intelligence requires active capability acquisition, yet current LLM agents inject pre-defined tool schemas into prompts, reducing models to passive selectors and falling short of robust general-purpose agency. We introduce MCP-Zero, an active agent framework that restores tool discovery autonomy to LLMs themselves. Instead of overwhelming models with all available tools, MCP-Zero enables agents to actively identify capability gaps, and request specific tools on-demand, transforming them from large-scale retrievers into genuine autonomous agents. The framework operates through three core mechanisms: (1) Active Tool Request, where models autonomously generate structured requests specifying their exact tool requirements; (2) Hierarchical Semantic Routing, a two-stage algorithm that matches requests to relevant servers and tools through improved semantic alignment; (3) Iterative Capability Extension, enabling agents to progressively build cross-domain toolchains while maintaining minimal context footprint. We construct MCP-tools, a comprehensive dataset of 308 MCP servers and 2,797 tools from the official Model-Context-Protocol repository. Experiments demonstrate that MCP-Zero preserves agent autonomy while achieving substantial efficiency gains: (i) accurate tool selection from nearly 3k candidates across 248.1k tokens; (ii) 98% reduction in token consumption on APIBank while maintaining high accuracy; and (iii) consistent multi-turn performance that scales with tool ecosystem growth. This work establishes active tool discovery as a fundamental design pattern for scalable autonomous agent systems. The code and dataset is released at https://github.com/xfey/MCP-Zero.
@@ 1 Introduction
We only think when we are confronted with a problem.
John Dewey
True thinking emerges not from passive consumption but from active engagement with challenges. Intelligence manifests through the capacity to actively seek solutions rather than merely selecting from predetermined options.
Large language models (LLMs) have evolved rapidly, demonstrating increasingly sophisticated reasoning and problem-solving capabilities [30, 8, 1]. The emergence of function-calling mechanisms has enabled large language models (LLMs) to transcend their parametric boundaries, leveraging external tools, APIs, and execution environments to accomplish complex real-world tasks [21, 26, 14, 18].
@@ 4.1 Dataset Construction
{mdframed}
[linewidth=1pt, linecolor=gray!50, backgroundcolor=gray!5]
@@
Figure 3: The schema of the MCP-tools dataset.
Structured Information Extraction. We employed Qwen2.5-72B-Instruct [24] with carefully designed few-shot examples to extract structured information following a standardized schema, which is demonstrated in Figure 3. The server_summary field represents a key innovation: LLM-generated comprehensive summaries that distill server capabilities while excluding operational details. This design enables effective semantic matching across diverse query formulations, from formal API specifications to natural language capability descriptions.